# Secure Control of Connected and Automated Vehicles Using Trust-Aware Robust Event-Triggered Control Barrier Functions

### 05 Jan 2024 v2

@@ IV-D Event-triggered Control
A common way to solve this dynamic optimization problem is to discretize $[t_{i}^{0},t_{i}^{f}]$ into intervals $[t_{i}^{0},t_{i}^{0}+\Delta],...,[t_{i}^{0}+k\Delta,t_{i}^{0}+(k+1)\Delta],...$ with equal length $\Delta$ and solving (16) over each time interval. The decision variables $u_{i,k}=u_{i}(t_{i,k})$ and $e_{i,k}=e_{i}(t_{i,k})$ are assumed to be constant on each interval and can be easily calculated at time $t_{i,k}=t_{i}^{0}+k\Delta$ through solving a QP at each time step:
@@
The set $\hat{C}_{i}\subset C_{i}$. because $b_{q}(\boldsymbol{\hat{x}}_{i,j}-\boldsymbol{w}_{i,j})\geq\min_{\{\boldsymbol{w}_{i,j}:\|\boldsymbol{w}_{i,j}\|_{\infty}\leq\epsilon_{1}\}}b_{q}(\boldsymbol{\hat{x}}_{i}-w_{i})\geq 0$. The minimum can be derived in closed form as shown below:
$\min_{\{\boldsymbol{w}_{i,j}:|\boldsymbol{w}_{i,j}\|_{\infty}\leq\epsilon_{1}\}}b_{q}(\boldsymbol{\hat{x}}_{i,j}-\boldsymbol{w}_{i,j})=\min_{\{\boldsymbol{w}_{i,j}:|\boldsymbol{w}_{i,j}\|_{\infty}\leq\epsilon_{1}\}}\hat{x}_{j}+w_{j}^{(x)}-\hat{x}_{i}-w_{i}^{(x)}-\varphi v_{i}-\Delta=\hat{x}_{j}-\hat{x}_{i}-\varphi v_{i}-\Delta-2\epsilon_{1}$ (22)
$\min_{\{\boldsymbol{w}_{i}:|\boldsymbol{w}_{i,j}\|_{\infty}\leq\epsilon_{1}\}}b_{3}(\boldsymbol{\hat{x}}_{i}-\boldsymbol{w}_{i})=\min_{\{\boldsymbol{w}_{i}:|\boldsymbol{w}_{i,j}\|_{\infty}\leq\epsilon_{1}\}}\hat{v}_{i}+w_{i}^{(v)}-v_{min}=\hat{v}_{i}-v_{min}-\epsilon_{1}$ (23)
$\min_{\{\boldsymbol{w}_{i}:|\boldsymbol{w}_{i,j}\|_{\infty}\leq\epsilon_{1}\}}\hat{v}_{max}-v-w_{i}^{(v)}=v_{max}-\hat{v}_{i}-\epsilon_{1}$ (24)
We can similarly define $\hat{X}_{j}(t_{i,k})$.
@@
where, $\boldsymbol{\hat{y}}_{i}=\boldsymbol{{y}}_{i}+\boldsymbol{w}_{i}$. The set $\hat{X}_{i}(t_{i,k})\subset X_{i}(t_{i,k})$. This is because we have the following
@@
Thus, $|\boldsymbol{\hat{y}}_{i}-\boldsymbol{\hat{x}}_{i}(t_{i,k})|\leq\boldsymbol{s}_{\boldsymbol{x}_{i}}-2[\epsilon_{1},\epsilon_{1}]^{T}\Rightarrow|\boldsymbol{y}_{i}-\boldsymbol{x}_{i}(t_{i,k})|\leq\boldsymbol{s}_{\boldsymbol{x}_{i}}$ Next, we seek a bound and a control law that satisfies the safety constraints within this bound. This can be accomplished by considering the minimum value of each component of (15) as shown next
@@
for $\forall q$. Note that the minimizations in (26), (28) and (27) are simple linear programs whose closed form solution can be derived. In order to apply this condition to the QP (17), we just replace (15) by (29) as follows:
@@
Finally, we can find $t_{i,k+1}$, time at which the first event(s) occur(s) since $t_{i,k}$ as a result of which the QP in (30) has to be solved as below:
@@
The following theorem formalizes our analysis by showing that if new constraints of the general form (29) hold, then our original CBF constraints (12) also hold. The proof follows the same lines as that of a more general theorem in [34] and, therefore, is omitted.
@@
The trust-based search guarantees safe coordination against Sybil attacks as proved in [23] and in conjunction with robust trust-aware CBFs from Corollary (1) makes our control and coordination framework safe against Sybil and Stealthy attacks. ∎
@@ V Attack Detection and Mitigation
Our proposed robust control scheme offers provably safe coordination against adversarial attacks. However, there are scenarios where attackers may target the network performance by causing traffic holdup. This is possible with Sybil attacks, as illustrated in [23], necessitating attack mitigation besides safety guarantee. The problem of detection involves the identification of adversarial (or, spoofed) CAVs accurately and mitigation can be defined as reestablishing the normal cooperation in the network close to what it would be in the ideal scenario without any attack. Resilience is necessary to ensure safe coordination until the attack is detected and in the presence of any false identification of adversarial (or, spoofed) CAVs. In this section, we present our novel mitigation framework based on the trust framework with the aforementioned objective.
@@ V-A Determination of Fake CAVs
Initially, every CAV is considered untrustworthy (i.e. $\tau_{i}(t_{i}^{0})=0$). Upon arrival in the CZ, the coordinator monitors the trust for each CAV and, if it detects any CAV $i\in S(t)$ s.t. $\tau_{i}(t)\leq 1-\delta$ and $\tau_{i}(t)\leq\tau_{i}(t-1)$, it initiates an observation window for that particular CAV of length $\eta$. If the trust for CAV $i$ is non-increasing and stays below the threshold of $1-\delta$ during the observation window then the coordinator proceeds to the mitigation step.
@@ V-B Robust Mitigation
The most trivial strategy that can be adopted is to rescind cooperation with the fake CAVs; however, it is essential to note that our presented framework can output false positives (although highly unlikely if the priorities of the behavioral specifications are chosen as mentioned in [20]) . Therefore, we offer a soft mitigation scheme; we call it soft because it is a passive scheme that relies on the local sensory information of the CAVs. This will become apparent in the remainder of the section. We define a rescheduling zone in the CZ of length $L_{1}$ as shown in Fig. 1. It has been shown that any passing sequence can be rescheduled in this area in [5]. Then, we present the following definitions.
@@
We mitigate the effect of fake CAVs by un-constraining the CAVs that are explicitly constrained by them (including the physically following CAVs if they are within their perception range and do not actually see any vehicle ahead) by solving the Integral Linear Program (ILP) defined below. Let the set of the ordered indices of detected fake CAVs that we want to mitigate be denoted as $S_{f}(t)$. We define the index $k_{min}=\min S_{f}\left(t\right)$ as the index of the first (fake) CAV in the queue to re-sequence from and $S_{+}(k_{min})=\{k_{min},\dots,N(t)\}$. Then, the ILP is formulated as follows:
$\max_{i\in S_{f}(t)}$ $\sum_{i\ \in S_{f}(t)}{a_{i}}$
@@
Based on the above definitions we now outline the scenarios that are of importance to us and derive an approximate solution of the ILP for them.
@@
    No CAVs are constrained by CAVs in $S_{f}(t)$: In this case, the solution of (34) will reschedule the CAVs starting from index $k=\min S_{f}(t)$ in $S_{f}(t)$ by moving them at the end of the queue and move the remaining CAVs with original index $i\geq k\text{ and }i\notin S_{f}(t)$ ahead in the queue to fill their places in their current order. This process will be repeated for $\forall k\in S_{f}(t)$.
@@
    There are CAVs in $S_{f}(t)$ which physically precede another CAV in the CZ: First, let us consider CAV $k\in S_{f}(t)$ and $j$ is the index of physically immediately following CAV, and let $S_{j}^{c}(t)\subseteq S(t)$ be the set of CAVs explicitly and implicitly constrained by $j$. At first, the CAVs with indices between $k$ to $j-1$ are moved ahead in the queue by incrementing their index by 1, and, then we make $k\leftarrow j-1$ where $j-1>k$. The reason for moving $k$ down the queue up to $j-1$ is because $k$ can be a real CAV which has been falsely identified as a fake CAV. Finally, remove the $\{j-1,j\}\cup S_{j}^{c}(t)$ from the queue, rearrange the queue by incrementing the indices of the remaining CAVs appropriately in the queue, and, add $\{j-1,j\}\cup S_{j}^{c}(t)$ in the queue. Then, repeat the process for the remaining CAVs in $S_{f}(t)$. Also update $S_{f}(t)$ accordingly.
    The final step is done to move any CAVs $k>j-1$ that are not explicitly constrained, or implicitly constrained by the immediately preceding CAV of CAV $j+1$ ahead of CAV $j-1$ in the queue.
@@
Observe that, for CAV $k\in S_{f}(t)$, upon rescheduling, the index of its immediately following CAV will become $k+1$. Once within the sensing range of CAV $k+1$, if CAV $k$ is not visible, it changes its control in (17) by removing the CBF constraint corresponding to CAV $i$ to complete the overtake. The coordinator detects the overtake completion by checking the satisfaction of the inequality in (35), upon which it completes the final step of the problem in (34) by swapping the indices of CAV $k$ and $k+1$ with each other. This step is performed $\forall k\ \in S_{f}(t)$ and repeated by following the scenarios mentioned previously (i.e. solution of (34)) until all fake CAVs reach the end of the queue.
@@
For FP cases, notice that for any $(j-1)\in S_{f}(t)$, $j$ is the CAV physically preceding it and every CAV $j^{+}>j$ that is not explicitly or implicitly constrained by $j$ are scheduled ahead of them after the first iteration of the algorithm. There will be no further rescheduling for $j-1$ meaning there will no cars overtaking it in the same road.
@@
In the rescheduling zone, any real CAV $i\in S(t)\backslash S_{f}(t)$ only overtakes a CAV $k\in S_{f}(t)$ if it doesn’t observe $k$ through its local perception. Similarly, any CAV $i\in S(t)\backslash S_{f}(t)$ only ignores the CBF condition in its control and jumps ahead of a CAV in $S_{f}(t)$ in the intersection if it doesn’t observe it through its local vision. This makes our proposed mitigation scheme soft (or, passive) and guarantees safety for false positive cases i.e. real CAVs which have been misidentified as fake CAVs. ∎
The fake CAVs are removed from the coordinator in one of two ways, namely: (i) the attacker stops sending information about a fake CAV, and (ii) the fake CAV leaves the CZ.
Figure 2: The values of average travel time, average energy, and average fuel consumption for real CAVs for different proportions of fake CAVs over 5 runs with our proposed mitigation scheme.
@@ VI Simulation Results
where we used typical values for parameters $\omega_{1},\omega_{2},\omega_{3},r_{0},r_{1}$ and $r_{2}$ as reported in [29]. The simulation was done in Sumo and Carla, where we used Sumo to generate various traffic scenarios and Carla to validate and evaluate the performance of our proposed schemes. The code for the simulated scenarios can be found in this link.
Trust-aware CBFs: We present results of comparison between trust aware robust CBFs with ordinary CBFs using event-triggered control framework. The results are summarized in Table I, containing simulation results for 30 vehicles with a Poisson traffic arrival process whose rate was set to 400 vehicles/hour. In the ordinary CBF case, the class $\mathrm{K}$ function is set to be linear function of its argument $\kappa_{q}=\kappa^{\prime}_{q}.b_{q}(.)$ where $\kappa^{\prime}_{q}=0.1$. We can see the benefits of incorporating the trust metric into CBFs, as there is a mixture of low-trust and high-trust vehicles. As can be seen, integrating trust makes the CBFs less conservative reducing the average travel times of the CAVs in the network and increasing average acceleration, thus improving the throughput of the network. However, this causes an increase in average fuel consumption expended by the vehicles in the CZ.
@@
Bias Injection Attack In order to highlight the robustness of our scheme against stealthy attacks and noise/estimation uncertainties we simulated an attack scenario by combining Sybil attack with BI attack. We compare our framework against the non-robust framework proposed in [20] and the results are shown in fig 4. As can be seen, the attack violates constraint (2) as shown in the plot of the constraint value (top left) which becomes negative due to the attack. This results in safety violation resulting in collision as shown in the image (on the left). On the other hand our proposed framework ensure safe coordination as can be verified from the plot and the image (on the right).
Figure 4: Results illustrating the merit of our proposed robust trust-aware event-triggered control scheme. The result was generated by simulating an attack scenario combining BI attack with Sybil attack. As can be seen, the framework in [20] results in safety violation (left) which is prevented by our proposed robust trust-aware event- triggered control scheme. The images shown above are from Carla simulator.
@@
Mitigation: The ultimate goal of having mitigation in place is to avoid accidents and minimize the effects of attacks on the performance of the traffic network (i.e., average travel time, average energy consumption, and average fuel consumption). We present our empirical results in Fig. 2 by injecting different proportions of fake CAVs during the attack and for each scenario performing 5 runs whose average and standard deviation is shown in the plots. We considered the strategic attacker model presented in [23]. It is important to note the considered model assumes that the attacker has no access to the RSU. We varied the location of the spoofed CAVs, their initial states, and the proportion of spoofed CAVs across the runs. As can be seen, with our proposed mitigation scheme the average travel time was almost reduced to the same value as the scenario with no attack, thus validating the efficacy of the mitigation scheme in maintaining network performance. In addition, the average energy was also reduced to almost what it was without an attack. We also notice that the average fuel consumption improves with our proposed mitigation scheme.
@@
Additionally, we provide a simulation scenario from CARLA during a Sybil attack in figure 3. The two figures shows the network performance with and without our proposed mitigation scheme after 1 minutes of starting the simulation. As can be seen, the absence of mitigation causes traffic holdup which is eased with our proposed mitigation scheme.
@@
False positive case. As mentioned our choice of trust framework doesn’t result in false positive cases. However, as our proposed method is invariant to the specific choice of the trust framework, hence we conducted experiments to analyze scenarios when a real CAV gets falsely identified as spoofed due to a poorly chosen trust framework. We conducted our experiments for various degrees of accuracy of the onboard vision system. For each scenario, we ran 100 experiments and computed the percentage of safe scenarios which is plotted in fig. 5. The experiments were run under different traffic conditions by varying the location of the falsely identified CAV (as spoofed) at the intersection for various values of states for the preceding car(s). An experiment was deemed safe if there were no collisions between real CAVs upon triggering mitigation. Our experiments illustrate we can guarantee safety with $95\%-99.96\%$ accuracy when the accuracy of the onboard objection detection pipeline varies from $85\%-95\%$.
@@
Figure 5: Plot of the percentage of safe scenarios over 100 runs for different degrees of accuracy of the onboard vision system.

### 25 Mar 2024 v3

@@ IV-D Event-triggered Control
A common way to solve (16)) is to discretize $[t_{i}^{0},t_{i}^{f}]$ into intervals $[t_{i}^{0},t_{i}^{0}+\Delta],...,[t_{i}^{0}+k\Delta,t_{i}^{0}+(k+1)\Delta],...$ with equal length $\Delta$ and solving (16) over each time interval. The decision variables $u_{i,k}=u_{i}(t_{i,k})$ and $e_{i,k}=e_{i}(t_{i,k})$ are assumed to be constant on each interval and can be easily calculated at time $t_{i,k}=t_{i}^{0}+k\Delta$ through solving a QP at each time step:
@@
$\min_{\{\boldsymbol{w}_{i}\|\boldsymbol{w}_{i}\|_{\infty}\leq\epsilon_{1}\}}b_{q}(\boldsymbol{\hat{x}}_{i}-\boldsymbol{w}_{i})\geq 0\Bigl{\}}$ (21)
Note that $\hat{C}_{i}\subset C_{i}$ because $b_{q}(\boldsymbol{\hat{x}}_{i,j}-\boldsymbol{w}_{i,j})\geq\min_{\{\boldsymbol{w}_{i,j}:\|\boldsymbol{w}_{i,j}\|_{\infty}\leq\epsilon_{1}\}}b_{q}(\boldsymbol{\hat{x}}_{i}-w_{i})\geq 0$. The minimum can be derived in closed form as shown below:
$\min_{\{\boldsymbol{w}_{i,j}:|\boldsymbol{w}_{i,j}\|_{\infty}\leq\epsilon_{1}\}}b_{q}(\boldsymbol{\hat{x}}_{i,j}-\boldsymbol{w}_{i,j})$
$=\min_{\{\boldsymbol{w}_{i,j}:|\boldsymbol{w}_{i,j}\|_{\infty}\leq\epsilon_{1}\}}\hat{x}_{j}-w_{j}^{(x)}-\hat{x}_{i}+w_{i}^{(x)}-\varphi(\hat{v}_{i}-w_{i}^{(v)})-\Delta$
$=\hat{x}_{j}-\hat{x}_{i}-\varphi v_{i}-\Delta-\epsilon_{1}(2+\varphi)$ (22)
$\min_{\{\boldsymbol{w}_{i}:|\boldsymbol{w}_{i,j}\|_{\infty}\leq\epsilon_{1}\}}b_{3}(\boldsymbol{\hat{x}}_{i}-\boldsymbol{w}_{i})$
$=\min_{\{\boldsymbol{w}_{i}:|\boldsymbol{w}_{i,j}\|_{\infty}\leq\epsilon_{1}\}}\hat{v}_{i}-w_{i}^{(v)}-v_{min}$
$=\hat{v}_{i}-v_{min}-\epsilon_{1}$ (23)
$\min_{\{\boldsymbol{w}_{i}:|\boldsymbol{w}_{i,j}\|_{\infty}\leq\epsilon_{1}\}}v_{max}-\hat{v}_{i}+w_{i}^{(v)}=v_{max}-\hat{v}_{i}-\epsilon_{1}$ (24)
We can similarly define $\hat{X}_{j}(t_{i,k})$:
@@
where $\boldsymbol{\hat{y}}_{i}=\boldsymbol{{y}}_{i}+\boldsymbol{w}_{i}$. Note that $\hat{X}_{i}(t_{i,k})\subset X_{i}(t_{i,k})$ since
@@
Thus, $|\boldsymbol{\hat{y}}_{i}-\boldsymbol{\hat{x}}_{i}(t_{i,k})|\leq\boldsymbol{s}_{\boldsymbol{x}_{i}}-2[\epsilon_{1},\epsilon_{1}]^{T}\Rightarrow|\boldsymbol{y}_{i}-\boldsymbol{x}_{i}(t_{i,k})|\leq\boldsymbol{s}_{\boldsymbol{x}_{i}}$.
Next, we seek a bound and a control law that satisfies the safety constraints within this bound. This can be accomplished by considering the minimum value of each component of (15) as shown next. For the first term, let
@@
for $q\in\{1,\dots,4\}$. Note that the minimizations in (26), (28) and (27) are simple linear programs whose closed form solution can be easily derived. In order to apply this condition to the QP (17), we just replace (15) by (29) as follows:
@@
Finally, we can determine $t_{i,k+1}$, the next time that a solution of the QP (30) must be solved, as follows:
@@
The following theorem formalizes our analysis by showing that if new constraints of the general form (29) hold, then our original CBF constraints (12) also hold. The proof follows the same lines as that of a more general theorem in [32].
@@
The trust-based search guarantees safe coordination against Sybil attacks as proved in [19] (in Theorem 1) and in conjunction with robust trust-aware CBFs from Corollary 1 makes our control and coordination framework safe against Sybil and Stealthy attacks. ∎
@@ V Attack Detection and Mitigation
Our proposed robust control scheme offers provably safe coordination against adversarial attacks. However, there are scenarios where attackers may target the network performance by causing traffic holdup. This is possible with Sybil attacks, as illustrated in [18], necessitating attack mitigation besides safety guarantees. The problem of detection involves the identification of adversarial (or, spoofed) CAVs accurately and mitigation can be defined as reestablishing the normal cooperation in the network close to what it would be in the ideal scenario without any attack. Resilience is necessary to ensure safe coordination until the attack is detected and in the presence of any false identification of adversarial (or, spoofed) CAVs. In this section, we present our proposed mitigation framework based on the trust framework with the aforementioned objective.
@@ V-A Determination of Fake CAVs
Initially, every CAV is considered untrustworthy (i.e., $\tau_{i}(t_{i}^{0})=0$). Upon arrival in the CZ, the coordinator monitors the trust for each CAV and, if it detects any CAV $i\in S(t)$ s.t. $\tau_{i}(t)\leq 1-\delta$ and $\tau_{i}(t)\leq\tau_{i}(t-1)$, it initiates an observation window for that particular CAV of length $\eta$. If the trust for CAV $i$ is non-increasing and stays below the threshold of $1-\delta$ during the observation window then the coordinator proceeds to the mitigation step.
@@ V-B Robust Mitigation
The most trivial strategy that can be adopted is to rescind cooperation with the fake CAVs; however, it is essential to note that our framework can output false positives (although highly unlikely if the priorities of the behavioral specifications are chosen as mentioned in [19]). Therefore, we offer a soft mitigation scheme; we call it “soft” because it is a passive scheme that relies on the local sensory information of the CAVs. This will become apparent in the remainder of the section. We define a rescheduling zone in the CZ of length $L_{1}$ as shown in Fig. 1. It has been shown that any passing sequence can be rescheduled in this area in [31]. Then, we present the following definitions.
@@
We mitigate the effect of fake CAVs by unconstraining the CAVs that are explicitly constrained by them (including the physically following CAVs if they are within their perception range and do not actually see any vehicle ahead) by solving the Integral Linear Program (ILP) defined below. Let the set of the ordered indices of detected fake CAVs that we want to mitigate be denoted as $S_{f}(t)$. We define the index $k_{min}=\min S_{f}\left(t\right)$ as the index of the first (fake) CAV in the queue to re-sequence from and $S_{+}(k_{min})=\{k_{min},\dots,N(t)\}$. Then, the ILP is formulated as follows:
$\max_{i\in S_{f}(t)}$ $\sum_{i\ \in S_{f}(t)}{a_{i}}$ (32)
@@
Based on the above definitions we now outline the scenarios that are of importance to us and derive an approximate solution of (32) for them.
@@
    No CAVs are constrained by CAVs in $S_{f}(t)$: In this case, the solution of (32) will reschedule the CAVs starting from index $k=\min S_{f}(t)$ in $S_{f}(t)$ by moving them at the end of the queue and move the remaining CAVs with original index $i\geq k\text{ and }i\notin S_{f}(t)$ ahead in the queue to fill their places in their current order. This process will be repeated $\forall k\in S_{f}(t)$.
@@
    There are CAVs in $S_{f}(t)$ which physically precede another CAV in the CZ: First, let us consider CAV $k\in S_{f}(t)$ and $j$ is the index of physically immediately following CAV, and let $S_{j}^{c}(t)\subseteq S(t)$ be the set of CAVs explicitly and implicitly constrained by $j$. At first, the CAVs with indices between $k$ to $j-1$ are moved ahead in the queue by incrementing their index by 1, then, we set $k\leftarrow j-1$ where $j-1>k$. The reason for moving $k$ down the queue up to $j-1$ is because $k$ can be a real CAV which has been falsely identified as a fake CAV. Finally, remove $\{j-1,j\}\cup S_{j}^{c}(t)$ from the queue, rearrange the queue by incrementing the indices of the remaining CAVs appropriately in the queue, and add $\{j-1,j\}\cup S_{j}^{c}(t)$ in the queue. Then, repeat the process for the remaining CAVs in $S_{f}(t)$. Finally, update $S_{f}(t)$ accordingly. The final step is needed to move any CAVs $k>j-1$ that are not explicitly constrained, or implicitly constrained by the immediately preceding CAV of CAV $j+1$ ahead of CAV $j-1$ in the queue.
@@
Moreover, observe that, for CAV $k\in S_{f}(t)$, upon rescheduling, the index of its immediately following CAV will become $k+1$. Once within the sensing range of CAV $k+1$, if CAV $k$ is not visible, it changes its control in (17) by removing the CBF constraint corresponding to CAV $i$ to complete the overtake. The coordinator detects the overtake completion by checking the satisfaction of the inequality in (36), upon which it completes the final step of the (32) by swapping the indices of CAV $k$ and $k+1$ with each other. This step is performed $\forall k\ \in S_{f}(t)$ and repeated by following the scenarios mentioned previously (i.e., the solution of ILP) until all fake CAVs reach the end of the queue.
@@
For FP cases, notice that for any $(j-1)\in S_{f}(t)$, $j$ is the CAV physically preceding it and every CAV $j^{+}>j$ that is not explicitly or implicitly constrained by $j$ is scheduled ahead of them after the first iteration of the algorithm. There will be no further rescheduling for $j-1$, i.e., there will be no vehicles overtaking it in the same road.
@@
In the rescheduling zone, any real CAV $i\in S(t)\backslash S_{f}(t)$ only overtakes a CAV $k\in S_{f}(t)$ if it does not observe $k$ through its local perception. Similarly, any CAV $i\in S(t)\backslash S_{f}(t)$ only ignores the CBF condition in its control and jumps ahead of a CAV in $S_{f}(t)$ in the intersection if it does not observe it through its local vision. This makes our proposed mitigation scheme soft (or passive) and guarantees safety for false positive cases i.e., real CAVs which have been misidentified as fake CAVs. ∎
The fake CAVs are removed from the coordinator queue in one of two ways: (i) the attacker stops sending information about a fake CAV, and (ii) the fake CAV leaves the CZ.
Figure 2: Results illustrating the merit of our proposed robust trust-aware event-triggered control scheme. The result was generated by simulating an attack scenario combining BI attack with Sybil attack. As can be seen, the framework in [19] results in safety violation (left) which is prevented by our proposed robust trust-aware event- triggered control scheme. The images shown above are from CARLA simulations.
Figure 3: The values of average travel time, average energy, and average fuel consumption for real CAVs for different proportions of fake CAVs over 5 runs with and without our proposed mitigation scheme.
@@ VI Simulation Results
where we used typical values for parameters $\omega_{1},\omega_{2},\omega_{3},r_{0},r_{1}$ and $r_{2}$ as reported in [33]. The simulation was done in Sumo and Carla, where we used Sumo to generate various traffic scenarios and Carla to validate and evaluate the performance of our proposed schemes.
Trust-aware CBFs: We present results comparing trust aware robust CBFs with ordinary CBFs using the event-triggered control framework. The results are summarized in Table I, containing simulations for 30 vehicles with a Poisson traffic arrival process whose rate was set to 400 vehicles/hour. In the ordinary CBF case, the class $\mathrm{K}$ function is set to be linear in its argument: $\kappa_{q}=\kappa^{\prime}_{q}.b_{q}(.)$ where $\kappa^{\prime}_{q}=0.1$. $\alpha$ is similar to as defined in (6). We can see the benefits of incorporating the trust metric into CBFs, as there is a mixture of low-trust and high-trust vehicles. As can be seen, integrating trust makes the CBFs less conservative reducing the average travel times of the CAVs in the network and increasing average acceleration, thus improving the throughput of the network. Finally, we notice that this also improves the average fuel consumption of the vehicles in the network.
@@
Bias Injection Attack In order to highlight the robustness of our scheme against stealthy attacks and noise/estimation uncertainties we simulated an attack scenario by combining Sybil attack with BI attack. We compare our framework against the non-robust framework proposed in [19] and the results are shown in Fig. 2. As can be seen, the attack violates constraint (2) as shown in the plot of the constraint value (top left) which becomes negative due to the attack. This results in safety violation resulting in collision as shown in the image (on the left). On the other hand our proposed framework ensure safe coordination as can be verified from the plot and the image (on the right).
@@
Mitigation: The ultimate goal of having mitigation in place is to avoid accidents and minimize the effects of attacks on the performance of the traffic network (i.e., average travel time, average energy consumption, and average fuel consumption). We present our empirical results in Fig. 3 by injecting different proportions of fake CAVs during the attack and for each scenario performing 5 runs whose average and standard deviation are shown in the plots. We considered the strategic attacker model presented in [18]. It is important to note that this model assumes that the attacker has no access to the RSU. We varied the location of the spoofed CAVs, their initial states, and the proportion of spoofed CAVs across the runs. As can be seen, with our proposed mitigation scheme the average travel time was reduced to almost the same value as the scenario with no attack, thus validating the efficacy of the mitigation scheme in maintaining network performance. In addition, the average energy was also reduced to almost what it was without an attack. Moreover, we notice that the average fuel consumption improves with our proposed mitigation scheme.
@@
Additionally, we provide a simulation scenario from CARLA during a Sybil attack in Fig. 4. The two figures show the network performance with and without our proposed mitigation scheme after 1 min. of starting the simulation. As can be seen, the absence of mitigation causes traffic holdup which is eased with our proposed mitigation scheme.
@@
False positive case. As mentioned, our choice of trust framework does not result in false positive cases. However, as our proposed method is invariant to the specific choice of the trust framework, we conducted experiments to analyze scenarios when a real CAV gets falsely identified as spoofed due to a poorly chosen trust framework. We conducted our experiments for various degrees of accuracy of the onboard vision system. For each scenario, we ran 100 experiments and computed the percentage of safe scenarios, with results shown in Fig. 5. The experiments were run under different traffic conditions by varying the location of the falsely identified CAV (as spoofed) at the intersection for various values of states for the preceding vehicle(s). An experiment was deemed “safe” if there were no collisions between real CAVs upon triggering mitigation. Our experiments show that we can guarantee safety with $95\%-99.96\%$ accuracy when the accuracy of the onboard object detection pipeline varies from $85\%-95\%$.
@@
Figure 5: Percentage of safe scenarios over 100 runs for different degrees of accuracy of the onboard vision system.